OpenStudy (anonymous):

Divide -3x3 - 2x2 - x - 2 by x – 2?

OpenStudy (megankeeping):

You can use factor theorem to work out the remainder f(x)=-3x3-2x2-x-2 f(2)=-3(2)3-2(2)2-(2)-2 = -24-8-2-2 = -36 The remainder is -36 You can use long division:  That shows that the answer is -3x2-8x-17 with remainder -36